Esperantologio / Esperanto Studies 2 (2001), 4

A word from the Editor

Christer Kiselman

D ear reader:

It is a pleasure to present to you the second issue of Esperanto Studies! It contains three original studies and four reviews. Of the articles, there is one in each of the languages Esperanto, English, and French; of the reviews, there are two in Esperanto, and one each in English and German.

For the editorial policy of the journal, please see the first issue.

⋆

Heidi Goes carried out a detailed study of the Esperantists in Tanzania and Togo in her licentiate dissertation, presented in 1999 to the University of Gent – it comprises 156 pages! Her article in this issue presents the principal findings of her study.

David Pardue has studied religion and the attempts toward language unity in Brazil, where spiritism and Esperanto live in symbiosis, and the language serves as more than a proselytization tool. He offers us some of his insights in a carefully documented article.

Walter Żelazny, a sociologist who works both in France and Poland, presents to us his thoughts on the social and religious ideas associated with the phenomenon of so-called artificial languages, setting them in a broad cultural and sociological framework.

⋆

In the review section we want to draw your attention to four important recent studies.

Christopher Gledhill published a grammar of Esperanto, whose most notable characteristic is that it is based on an analysis of a corpus of the language. Liu Haitao reviews the book in a nine-page essay.

Erich-Dieter Krause published a magnificent dictionary in 1999. Ebbe Vilborg reviews thoroughly all its (great) virtues and (small) demerits.

Anett Heil received her Ph.D. from the University of Rostock in 1998. Her dissertation is a comparative study of grammar reduction in French-based Creole languages and planned languages. Liu Haitao reviews her work.

Finally, Sabine Fiedler, who is already known to the readers of Esperanto Studies for her contribution to the first issue, presented her habilitation thesis to the University of Leipzig and published a book based on it. The topic is phraseology in planned languages. Till Dahlenburg reviews the book.

⋆

I hope you will enjoy reading these articles as much as I did!

⋆

My sincere thanks go to the referees, who have painstakingly and selflessly contributed to the quality of this issue, sharing their insights with the authors.